# Tackling Graph Oversquashing by Global and Local Non-Dissipativity

### v1

@@ Abstract
A common problem in Message-Passing Neural Networks is oversquashing – the limited ability to facilitate effective information flow between distant nodes. Oversquashing is attributed to the exponential decay in information transmission as node distances increase. This paper introduces a novel perspective to address oversquashing, leveraging properties of global and local non-dissipativity, that enable the maintenance of a constant information flow rate. Namely, we present SWAN, a uniquely parameterized model GNN with antisymmetry both in space and weight domains, as a means to obtain non-dissipativity. Our theoretical analysis asserts that by achieving these properties, SWAN offers an enhanced ability to transmit information over extended distances. Empirical evaluations on synthetic and real-world benchmarks that emphasize long-range interactions validate the theoretical understanding of SWAN, and its ability to mitigate oversquashing.
Machine Learning, ICML, oversquashing, over-squashing, deep graph network, graph neural network, long range interactions, ordinary differential equations
@@ 1 Introduction
A critical issue that limits Message-Passing Neural Networks (MPNNs) (Gilmer et al., 2017), a class of GNNs, is the oversquashing problem (Alon & Yahav, 2021; Di Giovanni et al., 2023). In the oversquashing scenario, the capacity of MPNNs to transmit information between nodes exponentially decreases as their distance increases, which imposes challenges on modeling long-range interactions, which are often necessary for real-world tasks (Dwivedi et al., 2022). At the same time, it has been shown in Haber et al. (2017); Chen et al. (2018a) that neural networks can be interpreted as the discretization of ordinary differential equations (ODEs). Building on these observations and understandings, similar concepts were utilized to forge the field of differential-equations inspired GNNs (DE-GNNs), as shown in (Poli et al., 2019; Chamberlain et al., 2021b) and subsequent works. Through this view, it is possible to design GNNs with strong inductive biases, such as smoothness (Chamberlain et al., 2021b), energy preservation (Eliasof et al., 2021; Rusch et al., 2022), node-wise non-dissipativity (Gravina et al., 2023), and more.
@@
Figure 1: An illustration of the ability of Global and Local Non-Dissipativity (e) to propagate information to distant nodes, from the source (a) to the target (b). Other dynamics, such as diffusion (c) cannot achieve the desired behavior, while Local Non-Dissipativity (d) has only limited effect. Our SWAN implements (e).
In this paper, we are interested in addressing the oversquashing problem in a principled manner, accompanied by theoretical understanding through the prism of DE-GNNs. Current literature offers several approaches to mitigate oversquashing, such as adding a virtual global node (Gilmer et al., 2017; Cai et al., 2023), graph rewiring (Gasteiger et al., 2019; Topping et al., 2022; Karhadkar et al., 2023; Nguyen et al., 2023), as well as using graph transformers (Shi et al., 2021; Dwivedi & Bresson, 2021; Rampášek et al., 2022). Some of the methods above focus on providing ad-hoc mechanisms for the network to reduce oversquashing. Instead, in this paper, we are interested in theoretically understanding and improving the information propagation capacity of the network, thereby mitigating oversquashing. Specifically, we will show that it is possible to design GNNs with a constant flow of information, unlike typical diffusion-based methods. To this end, we take inspiration from the recent ADGN (Gravina et al., 2023), a non-dissipative GNN. In ADGN, it was shown that by incorporating an antisymmetric transformation to the learned channel-mixing weights, it is possible to obtain a locally, i.e., node-wise, non-dissipative behavior. Here, we propose SWAN (Space-Weight ANtisymmetry), a novel GNN model that is both globally (i.e., graph-wise) and locally (i.e., node-wise) non-dissipative, achieved by space and weight antisymmetric parameterization. To understand the behavior of SWAN and its effectiveness in mitigating oversquashing, we propose a global, i.e., graph-wise, analysis, and show that compared to ADGN (Gravina et al., 2023), our SWAN is both globally and locally non-dissipative. The immediate implication of this property is that SWAN is guaranteed to have a constant information flow rate, thereby mitigating oversquashing. Such a property is visualized in Figure 1, where SWAN shows improved capacities of propagating information across the graph, with respect to diffusion and local non-dissipative approaches. To complement our theoretical analysis, we experiment with several long-range benchmarks, from synthetic to real-world datasets and tasks.
Main Contributions. In this paper, we present the following contributions. (i) A novel graph perspective theoretical analysis of the stability and non-dissipativity of antisymmetric DE-GNNs, providing a general design principle for introducing non-dissipativity as an inductive bias in any DE-GNN model. (ii) We propose SWAN, a space and weight antisymmetric DE-GNN with a constant information flow rate, and an increased distant node interaction sensitivity. (iii) We experimentally verify our theoretical understanding of SWAN, on both synthetic and real-world datasets. Our experiments show score improvements of up to 117% in synthetic datasets, as well as competitive results on real-world benchmarks compared to existing methods, from MPNNs to other DE-GNNs, and graph transformers – highlighting the importance of global and local non-dissipative behavior offered by SWAN.
@@ 2.1 Oversquashing
Oversquashing (Alon & Yahav, 2021) refers to the shortcoming of a GNN when transferring information between distant nodes, and is a common problem in Message-Passing Neural Networks (MPNNs). This shortcoming typically increases in tandem with the distance between the nodes, which hampers the ability of MPNNs to model complex behaviors that require long-range interactions. Standard MPNNs update node states by considering neighborhood information. To allow a node to receive information from $k$-hops distant nodes, an MPNN must employ at least $k$ layers, otherwise, it will suffer from under-reaching, because the two nodes are too far to interact with each other. However, the stacking of multiple message-passing layers also causes each node to receive an exponentially growing amount of information, as multiple hops are considered. This exponential growth of information, combined with the finite and fixed number of channels (features), can lead to a potential loss of information. As a consequence, the long-range effectiveness of MPNN is reduced.
@@ 2.2 GNNs Inspired by Differential-Equations
Notation. We consider a graph ${\mathcal{G}=(\mathcal{V},\mathcal{E})}$ as a system of interacting entities, referred as nodes, where $\mathcal{V}$ is a set of $n$ nodes, and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is a set of $m$ edges. We define the neighborhood of the $u$-th node as the set of nodes directly attached to it, i.e., $\mathcal{N}_{u}=\{v|(v,u)\in\mathcal{E}\}$. The $u$-th node is associated with a (possibly time-dependent) hidden feature vector $\mathbf{x}_{u}(t)\in\mathbb{R}^{d}$ with $d$ features, which provides a representation of the node at time $t$ in the system. The term $\mathbf{X}(t)=[\mathbf{x}_{0}(t),\ldots,\mathbf{x}_{n-1}(t)]^{\top}$ is an $n\times d$ matrix that represents the nodes state (features) at time $t$.
DE-GNNs. The core idea of DE-GNN architectures is to view GNNs as the discretization of the following ODE:
@@
where $\mathbf{x}_{u}(t=0)={\bf x}^{(0)}$ and $s\left(\{\mathbf{x}_{v}(t)\}_{v\in\mathcal{N}_{u}};\mathcal{G}\right)$ is a spatial aggregation function that depends on the graph $\mathcal{G}$ and the node features $\mathbf{x}_{u}(t)$. Specifically, it is common to implement $s\left(\{\mathbf{x}_{v}(t)\}_{v\in\mathcal{N}_{u}};\mathcal{G}\right)$ with graph diffusion, combined with a channel mixing operator implemented by a multilayer perceptron (MLP). Some examples of such methods were proposed in (Chamberlain et al., 2021b; Eliasof et al., 2021; Thorpe et al., 2022; Choi et al., 2023), and others.
Antisymmetric Weights in DE-GNNs. Learnable antisymmetric weights were studied in ADGN (Gravina et al., 2023), whose model is summarized below for completeness:
${\frac{\partial\mathbf{x}_{u}(t)}{\partial t}}=\sigma\Bigl{(}(\mathbf{W}-\mathbf{W}^{\top})\mathbf{x}_{u}(t)+\Phi(\{\mathbf{x}_{j}\}_{j\in\mathcal{N}_{u}},\mathbf{V})\Bigr{)},$ (2)
where ${\bf W},\ {\bf V}\in\mathbb{R}^{d\times d}$ are learnable weights, $\Phi(\{\mathbf{x}_{j}\}_{j\in\mathcal{N}_{u}},\mathbf{V})$ is any permutation invariant neighborhood aggregation function, and $\sigma$ is a non-linear activation function. The main theoretical property of ADGN (Gravina et al., 2023) is that it allows a stable and non-dissipative propagation from a local, node-wise perspective.
@@ 2.3 Mathematical Background
In this work, we are interested in studying the stability and non-dissipativity propagation of information in DE-GNNs. Therefore, we follow the analysis techniques presented in Chang et al. (2019); Gravina et al. (2023) and focus on analyzing the sensitivity of an ODE solution with respect to its initial condition, i.e.,
@@
Here, we present an overview of the various outcomes that can arise from this analysis.
We follow the results and assumptions from (Chang et al., 2019; Gravina et al., 2023), and consider the Jacobian, $\mathbf{J}(t)$ in Equation (3) to not change significantly over time. In Appendix 7, we provide a discussion of the justification as well as numerical verification of our the assumption. Given that, we can apply results from autonomous differential equations (Ascher & Petzold, 1998) and solve Equation 3 analytically as follows:
@@
Instability is observed when $\max_{i=1,...,d}Re(\lambda_{i}(\mathbf{J}))>0$. In this scenario, a small perturbation of the initial condition (i.e., a perturbation on the input graph) would cause an exponential divergence in node representations.
@@
Non-dissipativity is obtained when $Re(\lambda_{i}(\mathbf{J}))=0$. In this scenario, the system is stable, and the magnitude of $\partial\mathbf{x}(t)/\partial\mathbf{x}(0)$ is constant over time, and the input graph information is effectively propagated through the successive transformations into the final nodes’ representations.
@@ 3 SWAN: Space-Weight Antisymmetric GNN
Figure 2: The difference among non-dissipative and dissipative behaviors. With global (i.e., graph-wise) and local (i.e., node-wise) non-dissipative behavior (a), information is propagated between any pair of nodes with a viable path in the graph. Therefore, such a behavior increases the long-range effectiveness of the model, with propagation reaching distant nodes. A model exhibiting local non-dissipative behavior (b) enhances only the long-term memory capacity of individual nodes. A model demonstrating dissipative behavior (c) exhibits a convergence of node features toward non-informative values.
@@
$(\displaystyle\frac{\partial\mathbf{x}_{u}(t)}{\partial t}=\sigma\Bigl{(}$$(\mathbf{W}-\mathbf{W}^{\top})\mathbf{x}_{u}(t)$
$+\Phi(\{\mathbf{x}_{j}\}_{j\in\mathcal{N}_{u}},\mathbf{V})$
@@
$\begin{split}\Phi&=\sum_{v\in\mathcal{N}_{u}}(\hat{{\bf A}}_{uv}+\hat{{\bf A}}_{vu})(\mathbf{V}-{\bf V}^{\top})\mathbf{x}_{v}(t),\\ \Psi&=\sum_{v\in\mathcal{N}_{u}}((\tilde{{\bf A}}_{uv}-\tilde{{\bf A}}_{vu})(\mathbf{Z}+\mathbf{Z}^{\top})\mathbf{x}_{v}(t),\end{split}$ (6)
where $\tilde{\mathbf{A}},\hat{\mathbf{A}}\in\mathbb{R}^{n\times n}$ are neighborhood aggregation matrices that can be either pre-defined or learned. In our experiments, we consider two instances of Equation 6. The first, where $\tilde{\mathbf{A}},\hat{\mathbf{A}}$ are pre-defined by the random walk and symmetrically normalized adjacency matrices, respectively. The second, allows $\tilde{\mathbf{A}},\hat{\mathbf{A}}$ to be learned, as we describe later in Appendix A. More importantly, as we show now, the blueprint of $\Psi$ and $\Phi$, which is described in Equation 6 stems from a theoretical analysis of SWAN, that shows its ability to be non-dissipative both locally and globally , therefore leading to a globally constant information flow, regardless of time, that is, the model’s depth. Lastly, we note that the general formulation of $\Phi$ and $\Psi$ provide a general design principle for introducing non-dissipativity as an inductive bias in any DE-GNNs.
@@ 3.1 Node-wise Analysis of SWAN
We reformulate Section 3 to consider the formulation of $\Phi$ and $\Psi$ as in Equation 6, reading:
@@
SWAN is locally non-dissipative. Following the sensitivity analysis introduced in Section 2.3, we show that SWAN is stable and non-dissipative from the node perspective, i.e., it is locally non-dissipative. Therefore, we study the Jacobian matrix of our node-wise reformulation of SWAN.
In this case, the Jacobian $\mathbf{J}(t)=\mathbf{M}_{1}\mathbf{M}_{2}$ of Section 3.1 is composed of :
@@
$\mathbf{M}_{2}$ $=(\mathbf{W}-\mathbf{W}^{\top})+(\hat{{\bf A}}_{uu}+\hat{{\bf A}}_{uu})(\mathbf{V}-{\bf V}^{\top}).$ (9)
Following results from (Chang et al., 2019; Gravina et al., 2023), only the eigenvalues of $\mathbf{M}_{2}$ determine the local stability and non-dissipativity of the system in Section 3 for the final behavior of the model. Specifically, if the real part of all the eigenvalues of ${\bf M}_{2}$ is zero, then stability and non-dissipativity are achieved. We note that this is indeed the case in our system, since the real part of the eigenvalues of anti-symmetric matrices is zero, and ${\bf M}_{2}$ is composed of a summation of two anti-symmetric matrices.

### v2

# On Oversquashing in Graph Neural Networks Through the Lens of Dynamical Systems

@@ Abstract
A common problem in Message-Passing Neural Networks is oversquashing – the limited ability to facilitate effective information flow between distant nodes. Oversquashing is attributed to the exponential decay in information transmission as node distances increase. This paper introduces a novel perspective to address oversquashing, leveraging dynamical systems properties of global and local non-dissipativity, that enable the maintenance of a constant information flow rate. We present SWAN, a uniquely parameterized GNN model with antisymmetry both in space and weight domains, as a means to obtain non-dissipativity. Our theoretical analysis asserts that by implementing these properties, SWAN offers an enhanced ability to transmit information over extended distances. Empirical evaluations on synthetic and real-world benchmarks that emphasize long-range interactions validate the theoretical understanding of SWAN, and its ability to mitigate oversquashing.
@@ 1 Introduction
A critical issue that limits Message-Passing Neural Networks (MPNNs) (Gilmer et al. 2017), a class of GNNs, is the oversquashing problem (Alon and Yahav 2021; Di Giovanni et al. 2023). In the oversquashing scenario, the capacity of MPNNs to transmit information between nodes exponentially decreases as their distance increases, which imposes challenges on modeling long-range interactions, which are often necessary for real-world tasks (Dwivedi et al. 2022). At the same time, it has been shown in (Haber et al. 2018; Chen et al. 2018) that neural networks can be interpreted as the discretization of ordinary differential equations (ODEs). Building on these observations and understandings, similar concepts were utilized to forge the field of differential-equations inspired GNNs (DE-GNNs), as shown in (Poli et al. 2019; Chamberlain et al. 2021b) and subsequent works. Through this view, it is possible to design GNNs with strong inductive biases, such as smoothness (Chamberlain et al. 2021b), energy preservation (Eliasof, Haber, and Treister 2021; Rusch et al. 2022), node-wise non-dissipativity (Gravina, Bacciu, and Gallicchio 2023), and more.
In this paper, we are interested in addressing the oversquashing problem in a principled manner, accompanied by theoretical understanding through the prism of DE-GNNs. Current literature offers several approaches to mitigate oversquashing, such as adding a virtual global node (Gilmer et al. 2017; Cai et al. 2023), graph rewiring (Gasteiger, Weißenberger, and Günnemann 2019; Topping et al. 2022; Karhadkar, Banerjee, and Montufar 2023), as well as using graph transformers (Dwivedi and Bresson 2021; Rampášek et al. 2022). Some of the methods above focus on providing ad-hoc mechanisms for the network to reduce oversquashing. Instead, in this paper, we are interested in theoretically understanding and improving the information propagation capacity of the network, thereby mitigating oversquashing. Specifically, we will show that it is possible to design GNNs with a constant flow of information, unlike typical diffusion-based methods. To this end, we take inspiration from the recent ADGN (Gravina, Bacciu, and Gallicchio 2023), a non-dissipative GNN. In ADGN, it was shown that by incorporating an antisymmetric transformation to the learned channel-mixing weights, it is possible to obtain a locally, i.e., node-wise, non-dissipative behavior. Here, we propose SWAN (Space-Weight ANtisymmetry), a novel GNN model that is both globally (i.e., graph-wise) and locally (i.e., node-wise) non-dissipative, achieved by space and weight antisymmetric parameterization. To understand the behavior of SWAN and its effectiveness in mitigating oversquashing, we propose a global, i.e., graph-wise, analysis, and show that compared to ADGN (Gravina, Bacciu, and Gallicchio 2023), our SWAN is both globally and locally non-dissipative. The immediate implication of this property is that SWAN is guaranteed to have a constant information flow rate, thereby mitigating oversquashing. Such a property is visualized in Figure 1, where SWAN shows improved capacities of propagating information across the graph, with respect to diffusion and local non-dissipative approaches. To complement our theoretical analysis, we experiment with several synthetic and real-world long-range benchmarks.
@@
Figure 1: An illustration of the ability of Global and Local Non-Dissipativity in SWAN (e) to propagate information to distant nodes, from source (a) to the target (b). Other dynamics, such as diffusion (c) cannot achieve this behavior, while Local Non-Dissipativity (d) offers a limited effect.
Main Contributions. We present the following contributions. (i) A novel graph perspective theoretical analysis of the stability and non-dissipativity of antisymmetric DE-GNNs, providing a general design principle for introducing non-dissipativity as an inductive bias in any DE-GNN model. (ii) We propose SWAN, a space and weight antisymmetric DE-GNN with a constant information flow rate, and an increased distant node interaction sensitivity. (iii) We experimentally verify our theoretical understanding of SWAN, on both synthetic and real-world datasets. Our experiments show score improvements of up to 117% in synthetic datasets, as well as competitive results on real-world benchmarks compared to existing methods, from MPNNs to other DE-GNNs, and graph transformers – highlighting the importance of global and local non-dissipative behavior offered by SWAN.
@@ 2.1 Oversquashing
Oversquashing (Alon and Yahav 2021) refers to the shortcoming of a GNN when transferring information between distant nodes, and is a common problem in Message-Passing Neural Networks (MPNNs). Because standard MPNNs update node states by aggregating neighborhood information, oversqashing is amplified as node distances increase, hampering the ability of MPNNs to model complex behaviors that require long-range interactions. Namely, to allow a node to receive information from $k$-hops distant nodes, an MPNN must employ at least $k$ layers, otherwise, it will suffer from under-reaching, because the two nodes are too far to interact with each other. However, the stacking of multiple message-passing layers also causes each node to receive an exponentially growing amount of information, as multiple hops are considered. This exponential growth of information, combined with the finite and fixed number of channels (features), can lead to loss of information and reduced long-range effectiveness.
@@ 2.2 GNNs Inspired by Differential-Equations
Notation. We consider a graph ${\mathcal{G}=(\mathcal{V},\mathcal{E})}$ as a system of interacting entities, referred to as nodes, where $\mathcal{V}$ is a set of $n$ nodes, and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is a set of $m$ edges. We define the neighborhood of the $u$-th node as the set of nodes directly attached to it, i.e., $\mathcal{N}_{u}=\{v|(v,u)\in\mathcal{E}\}$. The $u$-th node is associated with a (possibly time-dependent) hidden feature vector $\mathbf{x}_{u}(t)\in\mathbb{R}^{d}$ with $d$ features, which provides a representation of the node at time $t$ in the system. The term $\mathbf{X}(t)=[\mathbf{x}_{0}(t),\ldots,\mathbf{x}_{n-1}(t)]^{\top}$ is an $n\times d$ matrix that represents the nodes state (features) at time $t$.
DE-GNNs. The core idea of DE-GNNs is to view GNNs as the discretization of the following ODE:
@@
where $\mathbf{x}_{u}(t=0)={\bf x}^{(0)}$ and $s\left(\{\mathbf{x}_{v}(t)\}_{v\in\mathcal{N}_{u}};\mathcal{G}\right)$ is a spatial aggregation function that depends on the graph $\mathcal{G}$ and the node features $\mathbf{x}_{u}(t)$. Specifically, it is common to implement $s\left(\{\mathbf{x}_{v}(t)\}_{v\in\mathcal{N}_{u}};\mathcal{G}\right)$ with graph diffusion, combined with a channel mixing operator implemented by a multilayer perceptron (MLP). Some examples of such methods were proposed in (Chamberlain et al. 2021b; Eliasof, Haber, and Treister 2021; Choi et al. 2023), and others.
Antisymmetric Weights in DE-GNNs. Learned antisymmetric weights were studied in ADGN (Gravina, Bacciu, and Gallicchio 2023), summarized below for completeness:
${\frac{\partial\mathbf{x}_{u}(t)}{\partial t}}=\sigma\Bigl{(}(\mathbf{W}-\mathbf{W}^{\top})\mathbf{x}_{u}(t)+\Phi(\{\mathbf{x}_{v}\}_{v\in\mathcal{N}_{u}},\mathbf{V})\Bigr{)},$ (2)
where ${\bf W},\ {\bf V}\in\mathbb{R}^{d\times d}$ are learnable weights, $\Phi(\{\mathbf{x}_{v}\}_{v\in\mathcal{N}_{u}},\mathbf{V})$ is any permutation invariant neighborhood aggregation function, and $\sigma$ is an activation function. The main theoretical property of ADGN (Gravina, Bacciu, and Gallicchio 2023) is that it allows a stable and non-dissipative propagation from a local, node-wise perspective.
@@ 2.3 Mathematical Background
While Alon and Yahav (2021) did not mathematically define oversquashing, recent works (Topping et al. 2022; Di Giovanni et al. 2023) associate it with exponentially declining node embedding sensitivity. Building on this perspective, we connect oversquashing to non-dissipative dynamical systems, leading to our SWAN. Specifically, we are interested in studying the stability and non-dissipativity propagation of information in DE-GNNs. Therefore, we follow the analysis techniques presented in (Chang et al. 2019; Gravina, Bacciu, and Gallicchio 2023) and focus on analyzing the sensitivity of an ODE solution with respect to its initial condition, i.e.,
@@
We now present an overview of the various outcomes that can arise from this analysis. We follow the results and assumptions from (Chang et al. 2019; Gravina, Bacciu, and Gallicchio 2023), and consider the Jacobian, $\mathbf{J}(t)$ in Equation 3 to not change significantly over time. We assume this condition in Theorems 3.1 and 3.2. In Appendix I, we provide a discussion of the justification as well as numerical verification of our assumption. Given that, we can apply results from autonomous differential equations (Ascher and Petzold 1998) and solve Equation 3 analytically as follows:
@@
Instability is observed when $\max_{i=1,...,d}Re(\lambda_{i}(\mathbf{J}))>0$, i.e., a small perturbation of the initial condition would cause an exponential divergence in node representations.
@@
Non-dissipativity is obtained when $Re(\lambda_{i}(\mathbf{J}))=0$. Here, the system is stable, and the magnitude of $\partial\mathbf{x}(t)/\partial\mathbf{x}(0)$ is constant over time, and the input graph information is effectively propagated through the successive transformations into the final nodes’ representations, addressing oversquashing.
@@ 3 SWAN: Space-Weight Antisymmetric GNN
Figure 2: The difference between non-dissipative and dissipative behaviors. With global (i.e., graph-wise) and local (i.e., node-wise) non-dissipative behavior (a), information is propagated between any pair of nodes with a viable path in the graph. Therefore, such a behavior increases the long-range effectiveness of the model. A model exhibiting local non-dissipative behavior (b) enhances only the long-term memory capacity of individual nodes. A model demonstrating dissipative behavior (c) exhibits a convergence of node features toward non-informative values.
@@
$(\displaystyle\frac{\partial\mathbf{x}_{u}(t)}{\partial t}=\sigma\Bigl{(}$$(\mathbf{W}-\mathbf{W}^{\top})\mathbf{x}_{u}(t)+\Phi(\{\mathbf{x}_{j}\}_{j\in\mathcal{N}_{u}},\mathbf{V})$
@@
$\Phi$ $=\sum_{v\in\mathcal{N}_{u}}(\hat{{\bf A}}_{uv}+\hat{{\bf A}}_{vu})(\mathbf{V}-{\bf V}^{\top})\mathbf{x}_{v}(t),$ (6)
$\Psi$ $=\sum_{v\in\mathcal{N}_{u}}(\tilde{{\bf A}}_{uv}-\tilde{{\bf A}}_{vu})(\mathbf{Z}+\mathbf{Z}^{\top})\mathbf{x}_{v}(t),$ (7)
where $\tilde{\mathbf{A}},\hat{\mathbf{A}}\in\mathbb{R}^{n\times n}$ are neighborhood aggregation matrices that can be pre-defined or learned. In our experiments, we consider two instances of Equations 6 and 7. The first, where $\tilde{\mathbf{A}},\hat{\mathbf{A}}$ are pre-defined by the random walk and symmetrically normalized adjacency matrices, respectively. The second, learns $\tilde{\mathbf{A}},\hat{\mathbf{A}}$, as described in Appendix A. Notably, as shown below, the parametrization of $\Psi$ and $\Phi$, described in Equations 6 and 7, ensures that SWAN’s node- and graph-wise Jacobians have purely imaginary eigenvalues (Sections 3.1 and 3.2), which, together with the background in Section 2.3, shows SWAN’s ability to be non-dissipative both locally and globally, leading to a globally constant information flow regardless of time, i.e., the model’s depth. Lastly, we note that the general formulation of $\Phi$ and $\Psi$ provide a general design principle for introducing non-dissipativity as an inductive bias in any DE-GNN (see Appendix A.4).
@@ 3.1 Node-wise Analysis of SWAN
We reformulate Section 3 to consider the formulation of $\Phi$ and $\Psi$ as in Equations 6 and 7, reading:
@@
SWAN is locally non-dissipative. Using the sensitivity analysis in Section 2.3, we show that SWAN is stable and non-dissipative from a node perspective, i.e., it is locally non-dissipative. In this case, the Jacobian $\mathbf{J}(t)=\mathbf{M}_{1}\mathbf{M}_{2}$ of Section 3.1 is composed of :
@@
$\mathbf{M}_{2}$ $=(\mathbf{W}-\mathbf{W}^{\top})+(\hat{{\bf A}}_{uv}+\hat{{\bf A}}_{vu})(\mathbf{V}-{\bf V}^{\top}).$ (10)
Following results from (Chang et al. 2019; Gravina, Bacciu, and Gallicchio 2023), only the eigenvalues of $\mathbf{M}_{2}$ determine the local stability and non-dissipativity of the system in Section 3 for the final behavior of the model. Specifically, if the real part of all the eigenvalues of ${\bf M}_{2}$ is zero, then stability and non-dissipativity are achieved. We note that this is indeed the case in our system, since the real part of the eigenvalues of antisymmetric matrices is zero, and ${\bf M}_{2}$ is composed of a summation of two antisymmetric matrices.